Three portions of the same liquid are mixed in a container that prevents the exchange of heat with the environment. Portion A has a mass m and a temperature of 95.0 °C, portion B also has a mass m but a temperature of 85.0 °C, and portion C has a mass mC and a temperature of 39.0 °C. What must be the mass of portion C so that the final temperature Tf of the three-portion mixture is Tf = 53.0 °C? Express your answer in terms of m; for example, mc = 2.20 m.
Answer #1

Let C is the specific heat of liquid.

Apply heat lost bet portionas A and B = heat gained by portion C

m*C*(95 - 53) + m*C*(85 - 53) = mC*C*(53 - 39)

m*(95 - 53) + m*(85 - 53) = mC*(53 - 39)

42*m + 32*m = 14*mC

==> mC = 74*m/14

mC = 5.286*m <<<<<<<<---------Answer